# Solving the 2D Advection-Diffusion Equation using Fixed-Depth Symbolic Regression and Symbolic Differentiation without Expression Trees

### v1

@@ Abstract
This paper presents a novel method for solving the 2D advection-diffusion equation using fixed-depth symbolic regression and symbolic differentiation without expression trees. The method is applied to two cases with distinct initial and boundary conditions, demonstrating its accuracy and ability to find approximate solutions efficiently. This framework offers a promising, scalable solution for finding approximate solutions to differential equations, with the potential for future improvements in computational performance and applicability to more complex systems involving vector-valued objectives.
@@ 1.1 Symbolic Regression
  • Leaf Nodes: Any of the individual features $\vec{x}=\{x_{1},x_{2},\ldots,x_{N}\}$ and constant tokens that can potentially be optimized with non-linear optimization routines such as L-BFGS [5] or Levenberg-Marquadt [13] [15].
@@ 1.2 Symbolic Differentiation
Symbolic differentiation is a computer algebra approach to solving derivatives analytically [8] through techniques such as the chain, product, and quotient rules. Symbolic differentiation falls under the umbrella of “computer algebra” technology [26], which modern-day software such as Mathematica [11], SymPy [17], Maxima [16] and many others have made generally available. However, as remarked in [21], many symbolic differentiation software rely on creating tree data structures for the input formulae and the corresponding derivatives. While the tree data structure enables versatility, it is not ideal for symbolic regression, which requires minimal memory allocations and maximal speed for optimal performance [27].
@@
In this work, an approximate analytical equation for $T$ is sought such that 3, subject to two different $\mathcal{I}$ and $\vec{\mathcal{B}}$ is approximately satisfied.
@@ 2 Related Work
In [19], the Euler-Bernoulli and Poisson equations subject to boundary constraints are solved using the Bingo framework [23] and automatic differentiation with Pytorch. [20]. The authors of [19] do not consider initial condition constraints, and they remark that differentiation with Pytorch performed sub-optimally compared to direct function evaluation.
@@ 3 Methods
Expressions are generated using the framework for fixed-depth grammar algorithms developed in [7] for prefix/postfix notation expressions. The algorithms implemented therein, namely, Random Search, Monte-Carlo Tree Search, Particle Swarm Optimization, Genetic Programming, and Simulated Annealing, are extended here to support multi-threading such that each thread runs a separate version of the search algorithm, and a concurrent map (using Boost’s concurrent flat map API [2]) containing the fitted constants of each unique expression encountered is shared.
@@
Once an expression $f$ is generated, it is plugged into the differential equation to produce another expression which is then evaluated on the considered domain to compute the mean-squared error. Exact derivatives are computed symbolically using the Reverse Polish Notation method developed in [21] for postfix expressions and the analogous Polish Notation method developed herein for prefix expressions.
@@ 4 Experiments and Results
Inspecting equation 3, one may notice that simply setting $T=0$, or anything independent of $x$, $y$, and $t$, is an exact solution. To avoid this outcome, it is required that the derivative of $T$ with respect to each of the input variables is nonzero up to a threshold, which is set to $\mathbf{1/\sqrt{2}}$ for the experiments in this section. If a generated expression does not satisfy this constraint, the mean-squared error is artificially set to $\infty$ to discourage trivial expressions.
@@ 4.2.1 Results
The configuration that resulted in the lowest Mean Squared Error (MSE) in this instance was configuration #3. Although configuration #1 was initially tested, it remained trapped in a local minimum for an extended period, yielding an MSE of 24.1156. Configuration #2 exhibited improved performance with an MSE of 0.0365547 yet also converged to a local minimum. Anticipating that Configuration #3 could potentially offer a more favorable bias-variance tradeoff, it was tested and yielded an MSE = 0.0135128 before reaching an impasse.
@@
$T$ $=\mathcal{I}^{\mathrm{sech}(0)}-\left(\left(\frac{y}{\left(\frac{4^{\mathcal{I}}}{\exp(x)}\right)}\right)-\left(\frac{y}{\left(\frac{4^{\mathcal{I}}}{\exp(x)}\right)}\right)\mathrm{tanh}((0.2-(t-20)))\right)$ (9)
$T$ $=\mathcal{I}-\frac{y\exp(x)}{4^{\mathcal{I}}}\left(1-\tanh(20.2-t))\right)$ (10)
Three plots for $t=0.1,10.5,20$ are shown in Figure 1. Movies can be found here.
Figure 1: Equation 10 for $t=0.1$, $t=10.5$, and $t=20$. The movies can be found here. This $T(x,y,t)$ distribution yielded an MSE = 0.0135128.
@@ 4.3.1 Results
The configuration that resulted in the lowest Mean Squared Error (MSE) in this instance was configuration #2. Configuration #1 was tested first, but after some time, it remained trapped in a local minimum for an extended period, yielding an MSE of 0.784878. Configuration #2 yielded the best result, an MSE = 0.689072 before reaching an impasse. Configuration #3 was tested several times, yet also converged to a local minimum, with the best MSE achieved being 0.749346.
@@
Three plots for $t=0.1,3.4,20$ are shown in Figure 2. Movies can be found here.
Figure 2: Equation 10 for $t=0.1$, $t=3.4$, and $t=20$. The movies can be found here. This $T(x,y,t)$ distribution yielded an MSE = 0.689072.
@@ 5 Remarks and Conclusions
In comparing Case 1 and Case 2, the simpler Case 1 led to a lower mean squared error (MSE). Case 2’s more complex velocity profile $\vec{u}$ and periodic boundary conditions resulted in a sparser solution space and costlier expression evaluations. For both cases, different algorithms showed varying success, highlighting the need for ongoing refinement and adaptation of search strategies.
Nevertheless, this study demonstrated an encourageable approach to solving the 2D advection-diffusion equation for two cases of initial and boundary constraints by employing fixed-depth symbolic regression and symbolic differentiation without expression trees. The methods utilized proved very efficient due to the elimination of expression trees and constraining the search to fixed-depth prefix/postfix symbolic expressions. The results of the two test cases showcased the ability of this approach to handle the respective boundary and initial conditions, though they could be extended to include further constraints to constrict the space of solutions allowed by physics.
This work lays the foundation for further exploration in applying symbolic regression and differentiation to more complex differential equations and systems involving vector-valued objectives, such as the Navier-Stokes equations. Future work could explore alternative search algorithms to improve convergence rates and minimize local minimum entrapments within the framework developed here.
Lastly, one potentially promising direction could be to develop a method in which $|\mathcal{D}(\hat{f})|$ can be estimated without having to evaluate data. This extension could be implemented with neural networks trained on the utilized basis set, or even simply assigning a magnitude rank to each operator and analyzing the structure of the corresponding expression tree. If successful, this extension would be expected to drastically improve efficiency, potentially making it tractable to brute-force search through the entire expression space for large expression depths, irrespective of the mesh granularity.

### v2

@@ Abstract
This paper presents a novel method for solving the 2D advection-diffusion equation using fixed-depth symbolic regression and symbolic differentiation without expression trees. The method is applied to two cases with distinct velocity fields and initial- and boundary-condition constraints. This method offers a promising solution for finding approximate analytical solutions to differential equations, with the potential for application to more complex systems involving vector-valued objectives.
@@ 1.1 Symbolic Regression
  • Leaf Nodes: Any of the individual features $\vec{x}=\{x_{1},x_{2},\ldots,x_{N}\}$ and constant tokens that can potentially be optimized with non-linear optimization routines such as L-BFGS [5] or Levenberg-Marquardt [13] [16].
@@ 1.2 Symbolic Differentiation
Symbolic differentiation is a computer algebra approach to solving derivatives analytically [8] through techniques such as the chain, product, and quotient rules. Symbolic differentiation falls under the umbrella of “computer algebra” technology [27], which modern-day software such as Mathematica [11], SymPy [18], Maxima [17], Maple [15], and many others have made generally available. However, as remarked in [22], many symbolic differentiation software rely on creating tree or graph data structures for the input formulae and the corresponding derivatives. While these data structure enables versatility, they may not be ideal for symbolic regression, which requires minimal memory allocations and maximal speed for optimal performance [28].
@@
In this work, approximate analytical equations for $T$ are sought such that 3, subject to two different $\mathcal{I}$ and $\vec{\mathcal{B}}$ is approximately satisfied.
@@ 2 Related Work
In [20], the Euler-Bernoulli and Poisson equations subject to boundary constraints are solved using the Bingo framework [24] with automatic differentiation done in Pytorch. [21]. The authors of [20] do not consider initial condition constraints, and they remark that differentiation with Pytorch performed sub-optimally compared to direct function evaluation.
@@ 3 Methods
Expressions are generated using the framework for fixed-depth grammar algorithms developed in [7] for prefix/postfix notation expressions. The algorithms implemented therein, namely, Random Search, Monte-Carlo Tree Search, Particle Swarm Optimization, Genetic Programming, and Simulated Annealing, are extended here to support multi-threading such that each thread runs a separate version of the search algorithm. A concurrent map (using Boost’s concurrent flat map API [2]) containing the fitted constants of each unique expression encountered is shared.
@@
Once an expression $f$ is generated, it is plugged into the differential equation to produce another expression which is then evaluated on the considered domain to compute the residual. Exact derivatives are computed symbolically using the Reverse Polish Notation method developed in [22] for postfix expressions and the analogous Polish Notation method developed herein for prefix expressions.
@@ 4 Experiments and Results
Inspecting equation 3, one may notice that simply setting $T=0$, or anything independent of $x$, $y$, and $t$, is an exact solution. To avoid this outcome, it is initially required that the derivative of $T$ with respect to each of the input variables is nonzero up to a threshold, which is set to $\mathbf{1/\sqrt{2}}$ for the experiments in this section. If a generated expression does not satisfy this constraint, the mean-squared error is artificially set to $\infty$ to discourage trivial expressions.
@@ 4.2.1 Results
The configuration that resulted in the lowest Mean Squared Error (MSE) in this instance was configuration #3. Although configuration #1 was initially tested, it remained trapped in a local minimum for an extended period, yielding an MSE of 24.1156. Configuration #2 exhibited improved performance with an MSE of 0.0365547 yet also converged to a local minimum. Anticipating that Configuration #3 could potentially offer a more favorable bias-variance tradeoff, it was tested and yielded an MSE = 0.00379332 before reaching an impasse.
@@
$T$ $=\left(\mathcal{I}^{\tanh(\mathcal{I})^{\sqrt{t}}}\right)-\left(\operatorname{sech}\left(\mathcal{I}+\frac{t}{0.2\cdot y}\right)\cdot\operatorname{sech}\left(x+\left(y+2^{\mathcal{I}}\right)\right)\right)$ (9)
$T$ $=\left(\mathcal{I}^{\tanh(\mathcal{I})^{\sqrt{t}}}\right)-\left(\operatorname{sech}\left(\mathcal{I}+\frac{t}{0.2\cdot y}\right)\cdot\operatorname{sech}\left(x+y+2^{\mathcal{I}}\right)\right)$ (10)
@@ 4.3.1 Results
The configuration that resulted in the lowest Mean Squared Error (MSE) in this instance was configuration #2. Configuration #1 was tested first, but after some time, it remained trapped in a local minimum for an extended period, yielding an MSE of 0.784878. Configuration #2 yielded the best result, an MSE = 0.689072 before reaching an impasse. Configuration #3 was tested several times and converged to a local minimum, with the best MSE achieved being 0.749346.
@@
4.3.2 Improved Results
The relatively large mean-squared error achieved for Case 2 in section 4.3.1 compared to the mean-squared error achieved for Case 1 in section 4.2.1 prompted further investigation into this case. Upon inspecting the corresponding expression tree for equation 16, it was observed that the expression could have been represented using a shallower tree. Thus, a depth of 5 was used in this section instead of 7, and the search was run again using GP (based on passed tests that had shown this algorithm to work well in practice). This search resulted in an improved MSE = 0.0117411 and the following expression before 17 and after 18 simplification.
$T$ $=\left(\mathcal{I}^{1^{\arcsin\left(\sqrt{\mathcal{I}}\right)}}\right)+\operatorname{sech}\left(\frac{\left(\frac{2}{6.283185}+(x+y)\right)}{(t\cdot 20)\cdot(t+0)}\right)$ (17)
$T$ $=\mathcal{I}+\operatorname{sech}\left(\frac{1/\pi+x+y}{20t^{2}}\right)$ (18)
4.4 Simulated Annealing of Best Results
To improve the results obtained in sections 4.2.1 and 4.3.2, the best-obtained expressions, namely, equations 10 and 18 for cases 1 and 2 respectively, were used as the starting expressions for the Simulated Annealing algorithm. Additionally, Particle Swarm Optimization was also enabled in this section, and, rather courageously, the constant threshold was reduced from $\mathbf{1/\sqrt{2}}$ to $\mathbf{0}$. Finally, to increase confidence in the results, the resulting expressions were annealed on $500\times 500\times 500$ mesh grids (as opposed to the $10\times 10\times 10$ mesh grids used previously).
4.4.1 Case 1
Using equation 9 as the starting point of this procedure, a new expression was found which, to floating-point (4-byte) precision, yielded an MSE = $\mathbf{0}$ (previously $3.79332\times 10^{-3}$), the resulting expression of which is given in equations 19 (un-simplified) and 20 (simplified).
$T$ $=\left(\mathcal{I}^{\operatorname{sech}(y_{0})^{\arcsin(0.1)}}\right)-\left(\operatorname{sech}\left((y-x)-\frac{\operatorname{sech}(0.103287)}{0.1^{\mathcal{I}}}\right)\cdot\frac{\frac{t}{\sin(\operatorname{sech}(1.1))}}{t-0.1^{0.1+1.1}}\right)$ (19)
$T$ $=\left(\mathcal{I}^{\operatorname{sech}(y_{0})^{\arcsin(0.1)}}\right)-\left(\operatorname{sech}\left((y-x)-\frac{\operatorname{sech}(0.103287)}{0.1^{\mathcal{I}}}\right)\cdot\frac{\frac{t}{\sin(\operatorname{sech}(1.1))}}{t-0.1^{1.2}}\right)$ (20)
4.4.2 Case 2
Employing the procedure in this section, equation 17 was improved to yield an MSE = $\mathbf{1.19209\times 10^{-7}}$ (previously $1.17411\times 10^{-2}$), the resulting expression of which is given in equations 21 (un-simplified) and 22 (simplified).
$T$ $=\left(\mathcal{I}^{y_{0}^{\log\left(\frac{6.283185}{6.283185}\right)}}\right)+\operatorname{sech}\left(\frac{\arccos(0.819757)+(x+y)}{t\cdot 20.0\cdot(12.499170)^{2}}\right)$ (21)
$T$ $=\mathcal{I}+\operatorname{sech}\left(\frac{\arccos(0.819757)+(x+y)}{t\cdot 20.0\cdot(12.499170)^{2}}\right)$ (22)
@@ 5 Remarks and Conclusions
In comparing Case 1 and Case 2, the simpler Case 1 led to a lower mean squared error (MSE). Case 2’s more complex velocity profile $\vec{u}$ and periodic boundary conditions resulted in a sparser solution space and costlier expression evaluations. For both cases, different algorithms showed varying success.
This study demonstrated an encourageable approach to finding approximate solutions to the 2D advection-diffusion equation for two cases of initial and boundary constraints by employing fixed-depth symbolic regression and symbolic differentiation without expression trees. The methods utilized proved very efficient due to the elimination of expression trees and constraining the search to fixed-depth prefix/postfix symbolic expressions. The results of the two test cases showcased the ability of this approach to handle the respective boundary and initial conditions, though they could be extended to include further constraints to constrict the space of solutions allowed by physics.
This work lays the foundation for further exploration in applying symbolic regression and differentiation to more complex differential equations and systems involving vector-valued objectives, such as the Navier-Stokes equations. Future work could explore alternative search algorithms to improve convergence rates and minimize local minimum entrapments within this framework.
Lastly, one potentially promising direction could be to develop a completely mesh-independent method of finding solutions to dynamical systems. This venture would include employing symbolic simplification to the expression $\mathcal{D}(\hat{f})$. In this scenario, an exact solution would yield zero terms on the left hand side, and an approximate solution would yield a left-hand-side small in the number of terms and estimated magnitude.